### v1

@@ 4.2 Comparisons with the Latest Methods
Results on Replica. We also evaluate our method in large-scale indoor scenes. Numerical valuations in Tab. 2 show that we produce much lower reconstruction errors than handcrafted renderers for UDFs, and even for SDFs. Visual comparisons in Fig. 8 show that our prior can recover geometry with higher accuracy, sharper edges, and much less artifacts. For objects that we can only observe from one side, our method is able to reconstruct it as a single surface, as illustrated in Fig. 9, which justifies the unsigned distance character. Please see our supplementary materials for evaluations on each scene.
Figure 11: Visualization of real scans used in NeUDF. The righttop and rightbottom part of each image are enlarged details and rendering views, respectively.
@@
Results on Real Scans. We further compare our method with the latest UDF reconstruction method NeUDF [29] on real scans in Fig. 1 and Fig. 10. We shot $4$ video clips on $4$ scenes with thin surfaces. The comparisons show that these challenging cases make NeUDF struggle to recover extremely thin surfaces like egg shells, resulting in incomplete and discontinuous surfaces. Comparing to the SDF learned by NeuS, our reconstruction with UDF produces much sharper structures. Similarly, we also produce more accurate and smoother surfaces than NeUDF on the real scans used in NeUDF, as shown in Fig. 11. Please watch our video for more details.
4.3 Ablation Studies and Analysis
@@
Geometry Overfitting with Depth Supervision. We first report analysis on the capability of geometry reconstruction from depth images on a single shape, which highlights the performance of our prior over others without the performance of color modeling. We learn a UDF with our prior or other handcrafted renderers from multiple depth images. Visual comparison in Fig. 12 shows that handcrafted renderers do not recover geometry detail even in an overfitting experiment, while our learned prior can recover more accurate geometry than others.
Neighboring Size. We report the effect of window size in our volume rendering prior on DF3D and DTU datasets in Fig. 13 and Fig. 14(a). We train different volume rendering priors with different window sizes including $\{1,10,20,30,40,50\}$. With a small window, such as $1$ and $10$, the prior becomes very sensitive to unsigned distance changes, which produces holes on the surface. With larger window sizes, such as $50$, the prior produces artifacts on the boundary. We find that a window covering $30$ queries works well in our experiments.
@@
Shapes for Learning Priors. We explore the effect of the number of shapes used for learning priors and report the generalization results on DF3D and DTU datasets, as represented in Fig. 14(b). The shapes are randomly selected from ShapeNet and DF3D datsets. The prior learned from a single shape exhibits a severe underfitting on DF3D, while more than two shapes do not bring further improvements. We further show the simplicity and robustness of learning our prior from different sets of shapes in Tab. 3. Each set still contains one randomly sampled cloth and one shape from different ShapeNet classes, which does not affect our appealing performance. The reason is that we sample lots of rays from different view angles to provide adequate knowledge of transforming unsigned distances to densities, which covers almost all unobserved situations in volume rendering for the UDF inference during testing. Additionally, calculating GT SDFs from GT meshes for every sampled point is a time-consuming operation when learning priors, therefore we select two shapes for both efficiency and performance.
(a) Neighboring Size
(b) Number of Shapes
Figure 14: Ablation study on the neighboring size and number of shapes. The numerical results are averaged across all scenes in DTU and DF3D datasets.
Queries for Implicit Representations. We justify the superiority of using sampling interval along the ray as queries. We try to remove the interval or replace the interval using other alternative like coordinates. The degenerated results in the “Inputs” column in Tab. 4 indicate that the relative position represented by the interval generate better on unobserved scenes.
Supervisions for Learning Priors. We further replace the supervisions of learning priors from depth images to RGB images or RGBD images, as reported in the “Supervisions” column in Tab. 4. Training with only RGB supervisions does not converge while the RGBD supervision severely degenerates the performance due to the aliasing of the color net. This indicates that the color affects the generalization ability of the prior a lot and is not suitable for learning priors for UDF rendering.
Table 3: Choice of different training shapes
Car+Cloth1 Chair+Cloth2 Sofa+Cloth3 Airplane+Cloth4 Bed+Cloth5
[…]
Fine-tuning Priors. Instead of using fixed parameters in the learned prior, we fine-tune the parameters of $f_{w}$ using RGB images as supervision during testing, as reported in the “Inference” column in Tab. 4. We find that the optimization does not converge. This indicates that the prior has acquired sufficient generalization ability during training, requiring no further adjustments during testing.
Table 4: Ablation studies on prior variants. The numerical results are averaged across all objects in DF3D dataset.
@@
Variants Ours only UDF UDF+Coor RGB Sup. RGBD Sup. Fine-tune
CD-L1 1.71 1.87 2.92 Divergent 2.02 Divergent
Figure 15: Ablation study on learning priors with SDF.
Learning with SDF. We also try to use our method to learn a prior for SDF instead of UDF with the same setting. Comparison in Fig. 15 shows similar results between UDF and SDF priors. The superior results over NeuS demonstrates the effectiveness of our approach of learning volume rendering priors for both UDFs and SDFs.
5 Conclusion
@@
We introduce volume rendering priors for UDF inference from multi-view images through neural rendering. We show that using data-driven manner to learn the prior can recover more accurate geometry than handcrafted equations in differentiable renderers. We successfully learn a prior from depth images from few shapes using our novel neural network and learning scheme, and robustly generalize the learned prior for UDFs inference from RGB images. We find that observing various unsigned distance variations during training and being 3D aware are the key to a prior with unbiasness, robustness, and scalability. Our extensive experiments and analysis on widely used benchmarks and real images justify our claims and demonstrate superiority over the state-of-the-art methods.
Acknowledgements
Yu-Shen Liu is the corresponding author. This work was supported by National Key R&D Program of China (2022YFC3800600), the National Natural Science Foundation of China (62272263, 62072268), and in part by Tsinghua-Kuaishou Institute of Future Media Data. We thank Junsheng Zhou, Takeshi Noda, You Peng for the discussions and the support for real-captured data. We also thank the anonymous reviewers for their efforts and valuable feedback to improve our work.
Appendix 0.A Appendix Overview
The appendix consists of additional results and implementation details. Section 0.B provides implementation details, including metrics, network structures, progressive learning strategy and data preparation. Section 0.C provides additional experimental results. Section 0.D discusses the limitations and future works.
Appendix 0.B Implementation Details
0.B.1 Metrics
For DTU dataset and DF3D dataset, we use Chamfer Distance (CD) as the metric. For Replica dataset, we report CD, Normal Consistency (N.C.) and F1-score following previous works [55, 3]. We use the algorithm introduced in MeshUDF [17] to extract surfaces from unsigned distance fields. To evaluate the reconstructed meshes in Replica dataset, we use the ground truth trajectory of the ground truth depth maps to detect the vertices which are visible by at least one camera. Triangles which have no visible vertices, either due to not being in any of the view frustum or due to being occluded by other surfaces, are culled. The point cloud to be evaluated is sampled on the culled mesh with a density of 1 point per square centimetre. The definitions of the metrics are the same as in [55].
0.B.2 Network Structures
We implement our volume rendering priors network $f_{w}$ as a 6-layer MLP with 256 hidden units and skip connections. We set both learning rate and weight decay to $1\times 10^{-4}$ for $f_{w}$ and train the network for 100k iterations with a batch size of 512. Similar to previous work [47, 30], we model the UDF function $f_{u}$ as an 8-layer MLP with 256 hidden units and skip connections, and the color function $f_{c}$ as a 4-layer MLP with 256 hidden units. We apply Softplus as the activation function after the last layer of $f_{u}$ to preserve the output of $f_{u}$ non-negative. The learning rate and decay scheduler of $f_{u}$ and $f_{c}$ are the same as NeuS [47]. We noticed that due to the non-differentiability of the unsigned distance function at the zero-level set, the normals near the surface are ambiguous. Therefore, we remove the normals from the input of the color network, which is a similar observation in NeuralUDF [30]. Same as the prior works [36, 47], we apply positional encoding to the spatial location with 6 frequencies and to the view direction with 4 frequencies.
0.B.3 Progressive Learning Strategy
There is a single trainable parameter called standard deviation in NeuS [47]. The trainable standard deviation helps SDF network to better capture the coarse shape at the early training stage. It is reduced with more training steps so that the surface becomes more clear and sharper. We design a similar mechanism for progressive learning of UDF in this work. Different from the learnable parameter used in NeuS, our prior parameterized by a neural network is fixed during the UDF inference. Thus, we use two different sets of parameters of the neural network during UDF inference. Specifically, we save parameters of prior network $f_{u}$ two times during learning the prior, one in the middle, and the other at the very end. The weight decay of the optimizer is reduced during the learning process. We observe that in the case of insufficient training and large weight decay, the network $f_{u}$ tends to learn a conservative transformation from UDF to opaque density. It is insensitive to UDF change at zero-level set, therefore mapping UDFs to relative small opaque densities, leading to a smooth distribution of rendering weights. At the end of training with small weight decay, the network tends to map UDFs to opaque densities in a radical way, which will make rendering weights concentrated at the surface. During the UDF inference, we first load the first set of parameters as our prior. This will recover a UDF representing coarse and incomplete geometry. Then, we load the second set of parameters, which will refine the UDF to recover geometry with details and sharp edges.
0.B.4 Data Preparation
To further demonstrate the generalization capability of our volume rendering priors in complex scenes, we manually capture four sets of data from real-world scenarios: an eggshell, a McDonald’s chicken nuggets box, a paper boat greeting card, and a potted plant. We use an iPhone 14 Plus to record a 2-minute object centered video with the resolution of 720$\times$720 for each scene. The videos are then cut into images at 2 frames per second, leading to a set of 120 images for each scene. Subsequently, we utilize COLMAP [41] to estimate the camera parameters, adjust the interest area of the estimated point cloud, and then generate camera parameters for training. Please refer to the video materials for visualization details.
Appendix 0.C Additional Results
0.C.1 Comparisons with Baselines
We provide detailed numerical comparisons against baseline methods for each scene in all datasets, along with additional visual comparisons.
Table 5: Quantitative evaluation results of Chamfer Distance of each object in DTU[21] dataset. We further finetune our method using patch loss [31] from the original NeuralUDF [30]. Our methods still outperforms NeuralUDF under the same experimental conditions.
Methods 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean
NeAT[35] 1.58 1.18 1.08 0.47 1.10 0.82 0.84 1.28 1.09 0.75 0.76 0.80 0.38 0.56 0.55 0.88
COLMAP[41] 0.81 2.05 0.73 1.22 1.79 1.58 1.02 3.05 1.40 2.05 1.00 1.34 0.49 0.78 1.17 1.36
NeuS[47] 1.37 1.21 0.73 0.40 1.20 0.70 0.72 1.01 1.16 0.82 0.66 1.69 0.39 0.49 0.51 0.87
NeuralUDF[30] 1.51 1.55 1.16 0.49 1.30 0.69 0.96 1.82 1.18 1.02 0.65 1.82 0.47 0.65 0.84 1.07
NeUDF[29] 2.27 2.79 2.11 0.73 2.52 0.97 1.63 1.53 1.63 0.68 0.93 3.44 0.59 0.82 0.99 1.58
Ours 0.94 1.17 0.84 0.47 1.25 0.68 0.64 1.57 1.02 0.96 0.60 1.33 0.35 0.49 0.50 0.85
NeuralUDF (w/ft) 0.69 1.18 0.67 0.44 0.90 0.66 0.67 1.32 0.94 0.95 0.57 0.86 0.37 0.56 0.55 0.75
Ours (w/ft) 0.55 0.95 0.45 0.43 0.94 0.75 0.61 1.40 1.01 0.74 0.60 0.94 0.31 0.50 0.50 0.71
Table 6: Quantitative evaluation results of Chamfer Distance (×$\times 10^{-3}$) of each object in DeepFashion3D[64] dataset. Note that NeAT uses additional mask supervision.
Methods 30 92 117 133 164 204 300 320 448 522 591 598 Mean
NeAT[35] 2.07 1.92 2.39 1.92 1.48 2.46 2.21 2.24 1.65 3.09 1.84 1.86 2.10
COLMAP[41] 2.95 2.91 3.58 3.06 3.23 3.26 3.09 3.11 2.95 3.16 2.97 2.95 3.10
NeuS[47] 3.18 4.82 4.78 4.99 3.73 5.71 5.89 2.21 5.89 3.60 2.44 5.13 4.36
NeuralUDF[30] 1.92 2.08 2.36 1.57 1.33 4.11 2.47 1.50 1.63 2.49 2.16 2.16 2.15
NeUDF[29] 1.96 2.45 1.58 1.57 1.32 3.62 1.90 1.45 1.44 2.85 2.18 1.79 2.01
Ours 1.59 1.73 2.06 1.63 1.44 2.07 1.66 1.60 1.39 2.14 1.50 1.67 1.71
Table 7: Quantitative evaluation results of each scene in Replica[42] dataset.
Methods room0 room1 room2 office0
[…]
NeAT[35] 0.237 0.715 0.348 0.274 0.683 0.271 0.144 0.785 0.305 0.098 0.828 0.566
COLMAP[41] 0.207 0.595 0.481 0.229 0.426 0.390 0.268 0.378 0.358 0.166 0.335 0.377
NeuS[47] 0.036 0.930 0.833 0.044 0.917 0.725 0.121 0.840 0.543 0.028 0.906 0.823
NeuralUDF[30] 0.122 0.838 0.462 0.072 0.888 0.595 0.094 0.883 0.565 0.089 0.826 0.589
NeUDF[29] 0.182 0.831 0.498 0.252 0.804 0.238 0.202 0.921 0.334 0.242 0.808 0.358
Ours 0.034 0.927 0.847 0.037 0.913 0.844 0.047 0.896 0.784 0.041 0.885 0.790
Methods office1 office2 office3 office4
[…]
NeAT[35] 0.207 0.707 0.310 0.152 0.781 0.402 0.155 0.790 0.403 0.224 0.784 0.308
COLMAP[41] 0.350 0.421 0.425 0.191 0.455 0.465 0.254 0.502 0.497 0.247 0.537 0.479
NeuS[47] 0.098 0.819 0.508 0.082 0.874 0.657 0.047 0.909 0.778 0.073 0.898 0.683
NeuralUDF[30] 0.112 0.818 0.424 0.126 0.845 0.557 0.135 0.858 0.583 0.137 0.887 0.473
NeUDF[29] 0.457 0.560 0.056 0.156 0.777 0.365 0.291 0.747 0.375 0.459 0.822 0.291
Ours 0.052 0.860 0.700 0.053 0.900 0.755 0.041 0.913 0.811 0.036 0.917 0.858
Figure 17: More visual comparisons between our method and baselines in DF3D [64] dataset.
Figure 18: More visual comparisons between our method and baselines in Replica [42] dataset.
DTU. We mentioned that NeuralUDF [30] uses additional patch loss [31] to fine-tune the resulted meshes, which is not the primary contribution of the differentiable renderer or used by other methods. Hence, for fair comparison, we report the results of NeuralUDF without fine-tuning across all datasets in the original paper. Here we integrate our method with patch loss and report the comparisons in Fig. 16, including our method and NeuralUDF with or without patch loss fine-tune, respectively. The visualization results show that our method, when fine-tuned using patch loss, still outperforms NeuralUDF under the same experimental conditions. Tab. 5 provides detailed numerical comparisons with baselines of each object in DTU dataset.
DeepFashion3D. We report the detailed numerical comparisons of each object in DeepFashion3D dataset in Tab. 6 and provide more visual comparisons, as shown in Fig. 17. Our volume rendering priors are capable of perceiving the 3D structure, thus recovering accurate wrinkles on the clothes, leading to small reconstruction error.
Replica. We provide numerical results of each scene in Replica dataset, as shown in Tab. 7. Additional visual comparisons are displayed in Fig. 18. When faced with complex indoor scenes, UDF-based methods consistently show poor performances. NeUDF [29] struggles to recover complete and continuous surfaces, NeuralUDF [30] exhibits large areas with blurry artifacts, while NeAT [35] produces erroneous layered structures. Our approach significantly surpasses these baselines and reconstructs accurate and smooth surfaces, which demonstrates the superiority of our method.
Table 8: Resource consumption of each procedure
Methods Steps Time Memory
Training Priors 100k 1.5h 3237MB
NeuS 300k 6.9h 6439MB
Training UDFs 300k 7.2h 7861MB
0.C.2 Training Resources
We report training steps, training time and memory cost in training and generalizing priors on DTU dataset, and compare with NeuS as a baseline in Tab. 8. Training priors require few resources, while the resource consumption of training UDF is similar with NeuS.
Appendix 0.D Discussion
0.D.1 Limitations
We adopt a data-driven based strategy for learning volume rendering priors and generalize it to other scenes during training UDFs. Although our prior only needs to be trained once for generalizing to unseen scenes, it unavoidably brings additional costs of storage and computational resources.
Additionally, similar to the common issues faced by multi-view neural implicit reconstruction methods, our approach fails to reconstruct low-texture regions, such as the area on the white plant pot in Figure 19. This problem might be further addressed by incorporating monocular depth and normal priors.
Figure 19: Failure case.
0.D.2 Future Works
A future work would be exploring a more elegant way of progressively learning UDFs instead of loading two-stage checkpoints. Additionally, it would be interesting to incorporate our learnable volume rendering priors into the recent neural rendering pipeline 3DGS [24, 19] to improve its quality of rendering and 3D reconstruction.

### v2

@@ IV-B4 Results on Replica
We also evaluate our method in indoor scenes. Numerical valuations in Tab. II show that we produce much lower reconstruction errors than handcrafted renderers for UDFs. Visual comparisons in Fig. 15 show that our prior can recover geometry with sharper edges and much less artifacts. For objects that are only observed from one side, we are able to reconstruct it as a single surface, as illustrated in Fig. 18, which justifies the unsigned distance character.
Figure 20: Comparison of UDF and SDF mesh extraction of our method on thin structures.
IV-B5 Results on Insects
We further evaluate our method in reconstructing complex open surfaces using Insects dataset. Numerical comparisons (CD $\times 10^{-3}$) in Tab. IV demonstrate the superiority of our method. Visualization results in Fig. 16 show that existing UDF methods often produce incomplete surfaces with a mixture of single and double layer patches, while our method accurately reconstructs them. We categorize thin structures into two types: single-layer planes and thin cubes. Our method performs well on the former, such as the butterfly wings. The latter, however, are more challenging to reconstruct since their extremely small surface area may not be covered by the uneven and relatively sparse sampling during volume rendering. Furthermore, the inherent ambiguity of UDFs near the zero-level set limits the reconstruction quality of thin cubes under existing UDF mesh extraction algorithms, which is not the primary focus of our method. We provide a visual comparison between UDF and SDF reconstructions in Fig. 20, where the SDF result recovers the thin cube structures with inflated thickness, such as mantis legs, indicating that our learned UDFs indeed capture these details, but not in a complete manner. In future work, we plan to explore more advanced UDF extraction techniques [24, 55] to further improve the reconstruction quality of thin structures.
IV-B6 Results on Real Scans
@@
We further compare our method with NeUDF [40] and VRPrior [77] on real scans in Fig. LABEL:fig:teaser and Fig. 19. We shot $4$ video clips on $4$ scenes with thin surfaces, including McNuggets, eggshell, paper boat and potted flower. The comparisons in Fig. LABEL:fig:teaser show that these challenging cases make NeUDF struggle to recover extremely thin surfaces like egg shells, resulting in incomplete and discontinuous surfaces. We also produce more accurate and smoother surfaces than VRPrior, for example, the ship’s mast and sails, as well as the branches and flowers in the plants, as shown in Fig. 19.
Figure 21: Visualization of our method integrated with different Gaussian Splatting methods. Left: Error maps of reconstructed meshes. Right: Gaussian centers around the ground truth surfaces. We learn much tighter Gaussian representations by leveraging VRPrior to impose a surface loss.
TABLE V: Comparison of CD($\downarrow$) for refining Gaussian Splatting using our VRP. We also report the average time (minutes) and memory (GB) consumption across three datasets.
Datasets DF3D DTU Insects Time Mem
GaussianUDF[39] 1.60 0.68 4.34 48 1.2
GaussianUDF+Ours 1.45 0.65 3.54 59 2.1
QGS[81] 4.46 0.54 3.57 21 2.0
QGS+Ours 2.97 0.53 3.13 33 3.1
IV-C Extension to Gaussian Splatting Reconstruction
Our pretrained volume rendering prior can serve as a general surface refiner that provides dense queries around the UDF zero-level set to optimize the Gaussian distributions toward the true surfaces, as described in Sec. III-F. We validate our efficiency by integrating our strategy into both GaussianUDF [39] and QGS [81], and conducting experiments on DF3D, DTU, and Insects datasets. We report the average CD as well as training time and memory consumption in Tab. V. The resource costs during inference are identical to those of Gaussian Splatting, as the prior network is not used at inference. Our method consistently improves performance across different variants, particularly on open surface datasets, where the reconstructed surfaces are more sensitive to Gaussian distributions. We further visualize the error maps of reconstructed meshes and the distributions of Gaussian centers in Fig. 21. The results demonstrate that our method effectively optimizes the Gaussian distributions toward the true surfaces, leading to more accurate and faithful reconstructions.
Figure 22: Effect of our volume rendering priors when we apply it to different neural representations, including signed distance functions (represented by NeuS [62]) and occupancy (represented by UNISURF [52]).
IV-D Extension to Different Neural Representations
Our volume rendering prior is not limited to unsigned distances. It can serve as a general module that transforms various neural implicit representations into rendering weights for alpha blending under different training settings. In this section, we evaluate our volume rendering priors when using SDF [53] and occupancy [47] as neural implicit representations to justify the generalizability of our framework.
IV-D1 Signed Distance Function
To learn the weight function $f_{ws}$ for SDF inference, we first sample ground truth signed distances $f_{s_{gt}}^{h}$ from the ground truth meshes $\{S_{h}\}_{h=1}^{H}$ and then replace the UDF input $f_{u_{gt}}^{h}$ in Eq. (2) with SDF input $f_{s_{gt}}^{h}$. All other settings are same as those used for training $f_{w}$. During generalizing, we use our learned $f_{ws}$ instead of the differentiable renderer of NeuS [62] to estimate the signed distance fields $f_{s}$. We report the reconstruction metrics on DTU dataset in Tab. VI. The superiority in numerical results beyond NeuS demonstrate the effectiveness of our volume rendering priors when we apply the prior to signed distance functions. Visualization comparisons in Fig. 22 show that our volume rendering prior corrects the reconstruction artifacts obtained with NeuS such as the holes and uneven surfaces.
TABLE VI: Quantitative evaluations in terms of Chamfer Distance based on different neural representations in DTU dataset.
Methods 24 37 40 55 63 97 110 122 Mean
NeuS[62] 1.07 1.21 0.73 0.40 1.20 1.16 1.59 0.51 0.98
NeuS+Ours(SDF) 0.94 1.05 0.56 0.42 1.14 0.98 1.33 0.52 0.87
Unisurf[52] 1.73 1.42 1.59 0.45 1.30 1.13 1.43 0.71 1.22
Unisurf+Ours(Occ) 1.25 1.17 1.11 0.52 1.22 0.97 1.25 0.62 1.01
IV-D2 Occupancy
We adopt the similar way of learning SDFs and UDFs to learn the weight function $f_{wo}$ for occupancy inference. Note that the ground truth occupancies $f_{o_{gt}}^{h}$ are binary values, which will hinder the generalization ability of the priors if trained directly on such data. To address this issue, we use ground truth SDFs to approximate the real distribution of occupancy fields,
$f_{o_{gt}}^{h}=\Phi(f_{s_{gt}}^{h}),\Phi(x)=\frac{1}{1+e^{20x}}.$ (8)
During inference, we accumulate the transformed occupancies $f_{wo}(f_{o})$ as the weights for alpha blending, rather than directly accumulating occupancy values $f_{o}$ as in UNISURF [52]. We also report numerical comparisons in Tab. VI, where our results significantly outperform UNISURF, revealing the potential of our volume rendering priors in mitigating the bias from various differentiable renderers and neural implicit representations. Visual comparison in Fig. 22 further shows that our volume rendering priors enable UNISURF to learn complete and smooth surfaces without holes.
@@ IV-D2 Occupancy
TABLE VII: Ablation study on the choice of neighboring size.
Neighboring Size Datasets
1 10 20 30 40 50 DTU[27] DF3D[89]
1.01 2.26
0.87 2.21
0.85 1.71
0.89 1.99
0.85 1.65
0.80 1.59
0.83 1.73
0.89 1.97
0.92 2.12
TABLE VIII: Ablation study on the multi-resolution windows and hierarchical sampling strategies.
DTU[27] DF3D[89]
VRPrior [77] 0.85 1.71
+Multi-resolution windows 0.82 1.67
+Point sampling priors 0.81 1.62
+Uniform sampling 0.80 1.59
@@
Figure 25: Ablation study on the number of shapes.
TABLE IX: Ablation studies on prior variants. The numerical results are averaged across all objects in DF3D dataset.
@@
Variants Ours only UDF UDF+Coor RGB Sup. RGBD Sup. Finetune
CD-L1 1.59 1.81 2.93 $\times$ 2.05 $\times$
IV-E Ablation Studies and Analysis
IV-E1 Geometry Overfitting with Depth Supervision
We first report analysis on the capability of geometry reconstruction from depth images on a single shape, which highlights the performance of our prior over others without the influence of color modeling. We learn a UDF with our prior or other renderers from multiple depth images. Visual comparison in Fig. 23 shows that handcrafted renderers do not recover geometry detail even in an overfitting experiment, while our learned prior can recover more accurate geometry than others.
IV-E2 Neighboring Size
We report the effect of window size in our volume rendering prior on DF3D and DTU datasets in Tab. VII, where we train different volume rendering priors with different single or multi-resolution window sizes including $\{1,10,20,30,40,50\}$. Compared to single window size, we find that multi-scale window sizes exhibit better generalization performance. This is because the multi-resolution setting enables the prior network to aggregate UDF variation information from different neighboring sizes, leading to a more accurate opaque density estimation. We also provide visualization results using different single or multiple window sizes, as shown in Fig. 24. Small windows, such as $\{1,10,20\}$, make the prior sensitive to unsigned distance changes, while large-span windows, such as $\{1,30,50\}$, introduce ambiguity by blending information from vastly different resolutions. We find that window sizes covering $\{10,20,30\}$ queries work well in all of our experiments.
IV-E3 Shapes for Learning Priors
We explore the effect of the number of shapes used for learning priors and report the generalization results on DF3D and DTU datasets, as represented in Fig. 25. The shapes are randomly selected from ShapeNet and DF3D datasets. The prior learned from a single shape exhibits a severe underfitting on DF3D, while more than six shapes do not bring further improvements. The reason is that we sample lots of rays from different view angles to provide adequate knowledge of transforming unsigned distances to densities, which covers almost all unobserved situations in volume rendering for the UDF inference during testing. Additionally, calculating GT UDFs from GT meshes for every sampled point is a time-consuming operation when learning priors, therefore we select six shapes for both efficiency and performance.
IV-E4 Effectiveness of Sampling Strategies
We additionally report effects of our technical contributions over VRPrior [77], as listed in Tab. VIII. Building upon VRPrior, we progressively integrate each one of our modules to show the improvements of the reconstructed results on DTU and DeepFashion3D datasets. The first modification replaces the single window size in VRPrior with our multi-resolution windows. The second one introduces point sampling priors as an auxiliary network to refine the sampling near zero-level set during UDF inference. And the third one incorporates our uniform sampling strategy to promote an even distribution of upsampled points. The ablation results demonstrate the effectiveness of ours proposed modules. Although the numerical gains are limited by the NeuS backbone, our approach can be further enhanced through integration with other frameworks. For example, incorporating patch loss yields a notable performance improvement, as reported in Tab. III, while combining our strategy with Gaussian-based reconstruction methods leads to substantial additional improvements, as demonstrated in Tab. V.
IV-E5 Variations of training VRPrior
We discuss several possible variations in the training of our volume rendering prior. Queries for Implicit Representations. We justify the superiority of using additional sampling interval along the ray as queries. We try to remove the interval or replace the interval using other alternatives like coordinates. The degenerated results in the “Inputs” column in Tab. IX indicate that the relative position represented by the interval generate better on unobserved scenes. Supervisions for Learning Priors. We further replace the supervisions of learning priors from depth images to RGB images or RGBD images, as reported in the “Supervisions” column in Tab. IX. Training with only RGB supervisions does not converge while the RGBD supervision severely degenerates the performance due to the aliasing of the color net. This indicates that the color affects the generalization ability of the prior a lot and is not suitable for learning priors for UDF rendering. Fine-tuning Priors. Instead of using fixed parameters in the learned prior, we fine-tune the parameters of $f_{w}$ along with the optimization of $f_{u}$ during generalizing, as reported in the “Inference” column in Tab. IX. We find that the optimization does not converge. This indicates that the prior has acquired sufficient generalization ability during training, requiring no further adjustments during testing.
V Conclusion
@@
We introduce volume rendering priors for UDF inference from multi-view images through neural rendering. We show that using data-driven manner to learn the prior can recover more accurate geometry than handcrafted equations in differentiable renderers. We successfully learn a prior from depth images from few shapes using our novel neural network and learning scheme, and robustly generalize the learned prior for UDFs inference from RGB images. We find that observing various unsigned distance variations during training and being 3D aware are the key to a prior with unbiasedness, robustness, and scalability. To further facilitate unbiased hierarchical sampling in UDF learning, we introduce point sampling priors and uniform sampling term as novel schemes. Extensive experiments and analysis on widely used benchmarks and real scans justify our claims and demonstrate both superiority over the state-of-the-art methods and strong generalization across various neural implicit representations.